OpenStudy (unofficialllyy):

An engineer estimated the weight of a steel beam to be 300 pounds. The actual weight of the beam was 304 pounds. Find the absolute error and the percent error of the engineer's estimate. If necessary, round your answers to the nearest tenth.

OpenStudy (theloshua):

absolute error: 304 - 300 = 4 percent error: |accepted value - experimental value| \ accepted value x 100% (304 - 400) = (4/304)(100) (.01315)(100) = 1.32